# Many-Body Localization in the Age of Classical Computing

### v2

@@ IX.1.1 Quantum avalanches
The quantum avalanche scenario of [158] describes the following mechanism for thermalizing the MBL subsystem $A$ due to the interaction with ergodic grain $B$. The spin in subsystem $A$ closest to subsystem $B$ will get thermalized if the correction $h_{1}$ (75) is of the order of unity. When the spin thermalizes, it becomes a member of the ergodic bubble, increasing the level density of the bubble by a factor of $2$, $\rho_{B}\to 2\rho_{B}$. As a result, the denominator in Eq. (75) is decreased twice, showing that the enlarged bath is more potent to thermalize further spins. However, also the strength of the coupling is decreased as $e^{-k/\xi}$, where $k$ represents the distance of a a spin in subsystem $A$ to the ergodic grain $B$. This gives rise to a non-trivial competition between the decreasing interaction and the growing density of states. After $k$ iterations of the process, the equation describing the thermalization of $k$-th spin away from the bath has the form
@@
Suppose the localization length $\xi$ is larger than the threshold value $\xi^{*}=1/\ln 2$. In that case, the growth of the density of states of the enlarged ergodic bubble compensates for the exponential decrease of the norm of the interaction term $H^{(j)}_{AB}$ in (71). Hence, the thermalizing avalanche propagates through the subsystem $A$, leading to its thermalization, and the LIOMs become destabilized by the ergodic inclusion. The avalanches seeded by rare ergodic grains propagate through the MBL regions, leading to the transition to the ergodic phase. In contrast, if $\xi<\xi^{*}$, the avalanche does not propagate throughout the lattice and hence this theory predicts absence of thermalization.
In passing, we note that the avalanche scenario relies crucially on the possibility of balancing the exponential decrease of the interaction matrix element (associated with the exponential decay of the LIOMs) and the exponential decrease of the level spacing of the ergodic grain. The balance may be found provided that the system is 1D, as argued above. In contrast, the avalanche mechanism predicts no MBL phase for larger dimensions. Additionally, the avalanches could not propagate in a 1D system with a superexponential decay of LIOMs [209].
@@
The avalanche model can be directly investigated numerically. To that end, one first rewrites $H_{A}$ in terms of LIOMs $I_{\alpha}$ and neglects the dephasing terms involving more than a single LIOM operator. Subsequently, the ergodic bubble is assumed to be modeled by a random matrix $R$. These suppositions, together with the form of the interaction term (71), lead to the Hamiltonian (11). The latter, dubbed a "quantum sun model," hosts an ergodicity-breaking phase transition [390, 694]. The position of the transition in the quantum sun model can be accurately pin-pointed with standard ergodicity breaking indicators such as the average gap ratio $\overline{r}$. For the quantum sun model, the gap ratio $\overline{r}$ follows the behavior illustrated in Fig. 2(a), up to a very mild drift of the crossing point with system size $L$. This feature is noteworthy since the quantum sun model is a genuinely interacting many-body system, with Thouless time scaling exponentially with $L$ at the transition, as discussed in Sec. VI.1.2. This makes the quantum sun model an exciting test bed for various phenomena that remain unclear in numerical studies of the disordered XXZ spin chain. One example is the many-body mobility edge whose signatures have been observed numerically in the disordered XXZ spin chain [439, 683, 110, 90] despite the theoretical arguments against its existence [159]. Numerical calculations in the quantum sun model provide clear evidence of the many-body mobility edge [477]. It must be noted that the eigenstates of the quantum sun model (11) in its non-ergodic phase are Fock space localized, i.e., their fractal dimensions $D_{q}$ vanish in the non-ergodic phase at sufficiently large system size [739]. This is in contrast to the eigenstates in the MBL regime of disordered XXZ and $J_{1}$-$J_{2}$ models, as discussed in Sec. VI.2.1.
@@ IX.1.2 Phenomenological investigations of MBL transition
It has been established that the renormalization group approaches relying on the avalanche mechanism lead to a BKT  [242, 177] universality class for the MBL transition. The associated length scale grows exponentially near the transition (15), hindering studies of properties of a BKT transition in numerical or experimental investigations. The analysis of [430] indicates that the critical properties of such a transition, when analyzed in systems of size up to $L$, can be well captured assuming a power-law divergence of the correlation length (14), with exponent $\nu(L)$ depending on $L$. For the proposed BKT-like scaling (differing from the BKT scaling by the presence of non-analytic terms in the RG equations), [430] show that $\nu(L)$ grows indefinitely with $L$. However, $\nu(L)$ growth with system size is extremely slow, and results for disordered spin chains of size as large as $L=10^{6}$ may be well described by a power-law scaling with $\nu(L=10^{6})\approx 3$. Hence, distinguishing the BKT scaling or the BKT-like scaling from a power-law divergence is a complex task even within the RG schemes, which provide access to large system sizes due to their polynomial in $L$ complexity [178, 177]. Consequently, telling apart the BKT-like scaling and power-law scalings in numerical studies of disordered spin chains is a formidable task due to the exponential increase of the dimension of the many-body Hilbert space, which restricts the available system sizes to $L\lessapprox 20$.
@@ IX.1.3 Probing avalanches with planted thermal inclusions
The rare thermal bubbles emerging in regions of anomalously weak disorder are necessarily present in sufficiently large disordered many-body systems, even deep in the MBL regime. The ergodic bubbles amount to the structure of entanglement clusters at the crossover between ergodic and MBL regimes [273], and are characterized by an enhanced decay of correlation functions of local observables. Nevertheless, detecting such rare regions requires complex tools, for instance, based on machine learning approaches [632] or analysis of pairwise correlation functions [269]. Hence, the relevance of the quantum avalanche mechanism for the dynamics of disordered spin chains is most conveniently studied in systems where a weakly disordered region is introduced intentionally in a controlled fashion.
To manually plant a thermal bubble, we enforce the disordered landscape shown in Fig. 14. To that end, the on-site magnetic fields $h_{i}$ in the XXZ spin chain (7) are taken as uncorrelated random variables,
@@
The avalanche scenario assumes that the ergodic bubble grows, thermalizing the neighboring spins, which enhances its capabilities to thermalize further spins, as reflected by Eq. (76). This hypothesis may be contrasted with an opposite scenario in which the ergodic bubble induces non-trivial dynamics of the neighboring LIOMs, which, however, do not influence their neighbors’ thermalization process, i.e., they do not become part of the ergodic bubble. In that case, the impact of the ergodic bubble decreases exponentially with the distance $r$ and the MBL in the subsystem $A$ remains stable. In that case, a decay rate $\Gamma_{j}$ of LIOM $I_{j}$ (72) at a distance $r$ from the ergodic bubble can be estimated by the Fermi golden rule [134] as
@@
The latter equation is valid only on timescales shorter than $\rho_{B}^{-1}$. At longer times, the discreteness of the spectrum is resolved, and, typically, the LIOM $I_{j}$ does not decay.
Both the avalanche scenario and the above perturbative, non-thermalizing picture involve assumptions that may be not met by the dynamics of disordered spin chains. Instead, the disordered spin chains may realize a scenario intermediate between the two situations. For that reason [361] proposed to monitor the process of the thermalization of the MBL subsystem using the two-point correlation function
$g^{(2)}(i)=\frac{1}{L_{B}}\sum_{j\in B}\left(\braket{S^{z}_{i}S^{z}_{j}}-\braket{S^{z}_{i}}\braket{S^{z}_{j}}\right).$ (80)
Assuming that the initial state is |$\ket{\phi_{\mathrm{in},n_{B}}}$ (72) and that $W_{A}$ corresponds to the MBL regime, the correlation function $g^{(2)}(i)$ falls-off exponentially with a characteristic length $\xi_{d}$
$|g^{(2)}(i)|=c\exp\left(-\frac{r}{\xi_{d}(t)}\right)$ (81)
@@
Investigations of the dynamics of $g^{(2)}(i)$ in the disordered XXZ spin chain [634] confirm the expected trends. The thermalization distance $\xi_{d}(t)$ grows faster than logarithmically in time for disorder strengths $W_{A}$ corresponding both to the ergodic and the crossover regimes of the XXZ spin chain, i.e., for $W_{A}\lessapprox W^{*}(L)$. In this regime of disorder strengths, the avalanches seeded by the ergodic bubble $B$ propagate throughout the subsystem $A$, causing its thermalization. In contrast, in the MBL regime, for $W_{A}\gtrapprox W^{*}(L)$, the saturation of the growth of thermalization distance $\xi_{d}(t)$ at times proportional to $\rho_{B}^{-1}$ is observed when the system is initialized in the state |$\ket{\phi_{\mathrm{in},n_{B}}}$, i.e., in an eigenstate of the MBL subsystem. Notably, the stop of the propagation of avalanches is observed at disorder strengths comparable to $W^{*}(L)$, even though the planted thermal inclusion would typically be found only in much larger spin chains. For instance, [634] considers thermal inclusion of size $L_{B}=6$ and with disorder strength $W_{B}=0.5$, while the disorder in the MBL subsystem is $W_{A}=6$. Without planting them manually, the density of such rare regions may be approximated as $(W_{B}/W_{A})^{L_{B}}$, which is of the order of $10^{-6}$ for the chosen parameters.
@@
The results of [634] are consistent with the relevance of the avalanche scenario for the thermalization of MBL blocks by ergodic inclusions in the dynamics of disordered spin chains. There are, however, two main caveats. The tendency towards ergodicity with increased time and length scales is similar to that of the uniformly disordered spin chains. This prevents distinguishing the dynamics of the MBL phase from a very slow thermalization. Moreover, the presence of two lengthscales $L_{A}$ and $L_{B}$ makes the extrapolation of the results to the thermodynamic limit even more complicated than in the case of uniformly disordered XXZ spin chains. Nonetheless, the two-point correlation function (80) can be probed experimentally in many-body systems with planted thermal inclusions as demonstrated in [361].
@@
Another approach to the investigation of the effects of planted regions of weak disorder was taken in Ref. [478], which considers the dynamics of the autocorrelation function $C(t)$ [^35] Strictly speaking, Ref. [478] considers dynamics of the imbalance $I(t)$ from the initial Néel state. the XXZ spin chain with on-site potential drawn according to Eq. (78). The decay of $C(t)$ is strongly enhanced in the presence of thermal inclusion compared to the situation when $W_{A}=W_{B}$. Moreover, the results for $C(t)$ in the presence of the thermal inclusion collapse on a single curve when the time is rescaled according to $t\to te^{-aW}$ where $a=2.1$ is a constant. The rescaling shows a significant degree of uniformity of the dynamics across a broad regime of disorder strengths, suggesting thermalization of the system in the presence of thermal inclusion. However, the system considered in [478] is periodically driven by turning on and off the coupling $H_{AB}$ between the ergodic inclusion and MBL region. The same system, but with coupling $H_{AB}$ present at all times, shows a much weaker tendency towards thermalization [634]. While the coupling to thermal inclusion indeed increases the tendency towards thermalization, the rescaling $t\to te^{-aW}$ does not provide a satisfactory collapse of the results for $C(t)$ in a broad regime of disorder strengths in the system without the periodic driving. This suggests disparate dynamics in the autonomous system at small and large $W_{A}$, leaving the question about its thermalization in the $L,t\to\infty$ limit open.
@@
The results of the studies of disordered many-body systems with planted thermal inclusions are consistent with the predictions of the avalanche scenario. However, the approach of [361, 634], based on investigation of the thermalization length $\xi_{d}(t)$, asserts the mere consistence of the thermalization in the subsystem $A$ with the propagation of avalanches. Therefore, it is important to note, the results of [634] do not exclude the possibility that the thermalization of the subsystem $A$ occurs due to microscopic mechanism different from the avalanche scenario but with similar predictions for the behavior of $\xi_{d}(t)$. Whether such a mechanism could be formulated remains an open question. Nevertheless, for that reason, the problem of pin-pointing the microscopic mechanism responsible for the emergence of ergodicity in disordered spin chains remains a challenge for further investigations of the dynamics of disordered many-body systems.

### v3

@@ IX.1.1 Quantum avalanches
The quantum avalanche scenario of [165] describes the following mechanism for thermalizing the MBL subsystem $A$ due to the interaction with the ergodic grain $B$. The spin in subsystem $A$ closest to subsystem $B$ will get thermalized if the correction $h_{1}$ (75) is of the order of unity. When the spin thermalizes, it becomes a member of the ergodic bubble, increasing the level density of the bubble by a factor of $2$, $\rho_{B}\to 2\rho_{B}$. As a result, the denominator in Eq. (75) is decreased twice, showing that the enlarged bath is more potent to thermalize further spins. However, also the strength of the coupling is decreased as $e^{-k/\xi}$, where $k$ represents the distance of a spin in subsystem $A$ to the ergodic grain $B$. This gives rise to a non-trivial competition between the decreasing interaction and the growing density of states. After $k$ iterations of the process, the equation describing the thermalization of $k$-th spin away from the bath has the form
@@
Suppose the localization length $\xi$ is larger than the threshold value $\xi^{*}=1/\ln 2$. In that case, the growth of the density of states of the enlarged ergodic bubble compensates for the exponential decrease of the norm of the interaction term $H^{(j)}_{AB}$ in (71). Hence, the thermalizing avalanche propagates through the subsystem $A$, leading to its thermalization, and the LIOMs become destabilized by the ergodic inclusion. The avalanches seeded by rare ergodic grains propagate through the MBL regions, leading to the transition to the ergodic phase. In contrast, if $\xi<\xi^{*}$, the avalanche does not propagate throughout the lattice and hence this theory predicts absence of thermalization. A strong approximation involved in the latter statement is that the localization length $\xi$, introduced in the non-interacting picture, remains unchanged when adding interactions.
In passing, we note that the avalanche scenario relies crucially on the possibility of balancing the exponential decrease of the interaction matrix element (associated with the exponential decay of the LIOMs) and the exponential decrease of the level spacing of the ergodic grain. The balance may be found provided that the system is one dimensional, as argued above. In contrast, the avalanche mechanism predicts no MBL phase for higher dimensions. Additionally, the avalanches could not propagate in a 1D system with a superexponential decay of LIOMs [223].
@@
The avalanche model can be directly investigated numerically. To that end, one first rewrites $H_{A}$ in terms of LIOMs $\hat{I}_{\alpha}$ and neglects the dephasing terms involving more than a single LIOM operator. Subsequently, the ergodic bubble is assumed to be modeled by a random matrix $\hat{R}$. These suppositions, together with the form of the interaction term (71), lead to the Hamiltonian (11). The latter, dubbed a "quantum sun model", hosts an ergodicity-breaking phase transition [427, 750]. The position of the transition in the quantum sun model can be accurately pin-pointed with standard ergodicity breaking indicators such as the average gap ratio $\overline{r}$. For the quantum sun model, the gap ratio $\overline{r}$ follows the behavior illustrated in Fig. 2(a), up to a very mild drift of the crossing point with system size $L$. This feature is noteworthy since the quantum sun model is a genuinely interacting many-body system, with Thouless time scaling exponentially with $L$ at the transition, as discussed in Sec. VI.1.2. This makes the quantum sun model an exciting test bed for various phenomena that remain unclear in numerical studies of the disordered XXZ spin chains. One example is the many-body mobility edge whose signatures have been observed numerically in the disordered XXZ spin chain [477, 738, 114, 93] despite the theoretical arguments against its existence [166]. Numerical calculations in the quantum sun model provide clear evidence of the many-body mobility edge [515]. Another new development represents a detailed understanding of the fate of the observable matrix elements when the ergodicity breaking transition is approached from the ergodic side. In the quantum sun model, the ETH ansatz from Eq. (3) breaks down while the short-range spectral statistics still comply with the GOE predictions [356]. This phenomenon, dubbed fading ergodicity, provides a theoretical framework to detect proximity of the ergodicity breaking transition despite the system being ergodic and the observables thermalize after a quantum quench.
Recently, [744] argued that the corresponding RMT-like model description of the avalanche physics is the ultrametric model, defined on a Fock space. Carrying out a detailed comparison between the later model and the quantum sun model (11), the authors showed that both models exhibit a nearly identical value of the ergodicity breaking transition point. Moreover, building on the results for the ultrametric random matrices [229], it was argued that the eigenstates on the non-ergodic side of the quantum sun model are Fock-space localized, i.e., their fractal dimensions $D_{q}$ vanish in the non-ergodic phase at sufficiently large system size [744]. This is in contrast to the eigenstates in the MBL regime of the disordered XXZ and $J_{1}$-$J_{2}$ models, as discussed in Sec. VI.2.1.
The ultrametric model on a Fock space defines a general class of interacting models that describe the interaction between the pinned impurities and the remaining (non-interacting) degrees of freedom. Beside the avalanche models (e.g., the quantum sun model), another subset of models that share similar geometry are the central spin models, see, e.g., [531]. Future works should explore in more details the relevance of the ultrametric model on a Fock space for the physics of the disordered spin chains.
@@ IX.1.2 Phenomenological investigations of MBL transition
It has been established that the renormalization group approaches relying on the avalanche mechanism lead to a BKT universality class for the MBL transition [258, 187]. The associated length scale grows exponentially near the transition (15), hindering studies of properties of a BKT transition in numerical or experimental investigations. The analysis of [466] indicates that the critical properties of such a transition, when analyzed in systems of size up to $L$, can be well captured assuming a power-law divergence of the correlation length (14), with exponent $\nu(L)$ depending on $L$. For the proposed BKT-like scaling (differing from the BKT scaling by the presence of non-analytic terms in the RG equations), [466] showed that $\nu(L)$ grows indefinitely with $L$. However, $\nu(L)$ growth with system size is extremely slow, and results for the disordered spin chains of size as large as $L=10^{6}$ may be well described by a power-law scaling with $\nu(L=10^{6})\approx 3$. Hence, distinguishing the BKT scaling or the BKT-like scaling from a power-law divergence is a complex task even within the RG schemes, which provide access to large system sizes due to their polynomial-in-$L$ complexity [188, 187]. Consequently, telling apart the BKT-like scaling and power-law scalings in numerical studies of disordered spin chains is a formidable task due to the exponential increase of the dimension of the many-body Hilbert space, which restricts the available system sizes to $L\lessapprox 20$.
@@ IX.1.3 Probing avalanches with planted thermal inclusions
The rare thermal bubbles emerging in regions of anomalously weak disorder are necessarily present in sufficiently large disordered many-body systems, even deep in the MBL regime. The ergodic bubbles amount to the structure of entanglement clusters at the crossover between the ergodic and MBL regimes [292], and are characterized by an enhanced decay of correlation functions of local observables. Nevertheless, detecting such rare regions requires complex tools, for instance, based on machine learning approaches [683] or analysis of pairwise correlation functions [288]. Hence, the relevance of the quantum avalanche mechanism for the dynamics of disordered spin chains is most conveniently studied in systems where a weakly disordered region is introduced intentionally in a controlled fashion.
To manually plant a thermal bubble, we enforce the disordered landscape shown in Fig. 14. To that end, the on-site magnetic fields $h_{i}$ in the disordered XXZ spin chain (7) are taken as uncorrelated random variables,
@@
The avalanche scenario assumes that the ergodic bubble grows, thermalizing the neighboring spins, which enhances its capabilities to thermalize further spins, as reflected by Eq. (76). This hypothesis may be contrasted with an opposite scenario in which the ergodic bubble induces non-trivial dynamics of the neighboring LIOMs, which, however, do not influence their neighbors’ thermalization process, i.e., they do not become part of the ergodic bubble. In that case, the impact of the ergodic bubble decreases exponentially with the distance $r$ and the MBL in the subsystem $A$ remains stable. In that case, a decay rate $\Gamma_{j}$ of LIOM $\hat{I}_{j}$ (72) at a distance $r$ from the ergodic bubble can be estimated by the Fermi golden rule [139] as
@@
The latter equation is valid only on timescales shorter than $\rho_{B}^{-1}$. At longer times, the discreteness of the spectrum is resolved, and, typically, the LIOM $\hat{I}_{j}$ does not decay.
Both the avalanche scenario and the above perturbative, non-thermalizing picture involve assumptions that may not be met by the dynamics of the disordered spin chains. Instead, the disordered spin chains may realize a scenario intermediate between the two situations. For that reason, [393] proposed to monitor the process of thermalization of the MBL subsystem using the two-point correlation function,
$g^{(2)}(i)=\frac{1}{L_{B}}\sum_{j\in B}\left(\braket{\hat{S}^{z}_{i}\hat{S}^{z}_{j}}-\braket{\hat{S}^{z}_{i}}\braket{\hat{S}^{z}_{j}}\right).$ (80)
Assuming that the initial state is |$\ket{\phi_{\mathrm{in},n_{B}}}$ (72) and that $W_{A}$ corresponds to the MBL regime, the correlation function $g^{(2)}(i)$ falls-off exponentially with a characteristic length $\xi_{d}$,
$|g^{(2)}(i)|=c\exp\left(-\frac{r}{\xi_{d}(t)}\right),$ (81)
@@
Investigations of the dynamics of $g^{(2)}(i)$ in the disordered XXZ spin chain [685] confirm the expected trends. The thermalization distance $\xi_{d}(t)$ grows faster than logarithmically in time for disorder strengths $W_{A}$ corresponding both to the ergodic and the crossover regimes of the disordered XXZ spin chain, i.e., for $W_{A}\lessapprox W^{*}(L)$. In this regime of disorder strengths, the avalanches seeded by the ergodic bubble $B$ propagate throughout the subsystem $A$, causing its thermalization. In contrast, in the MBL regime, for $W_{A}\gtrapprox W^{*}(L)$, the saturation of the growth of thermalization distance $\xi_{d}(t)$ at times proportional to $\rho_{B}^{-1}$ is observed when the system is initialized in the state |$\ket{\phi_{\mathrm{in},n_{B}}}$, i.e., in an eigenstate of the MBL subsystem. Notably, the stop of the propagation of avalanches is observed at disorder strengths comparable to $W^{*}(L)$, even though the planted thermal inclusion would typically be found only in much larger spin chains. For instance, [685] considers thermal inclusion of size $L_{B}=6$ and with disorder strength $W_{B}=0.5$, while the disorder in the MBL subsystem is $W_{A}=6$. Without planting them manually, the density of such rare regions may be approximated as $(W_{B}/W_{A})^{L_{B}}$, which is of the order of $10^{-6}$ for the chosen parameters.
@@
The results of [685] are consistent with the relevance of the avalanche scenario for the thermalization of MBL blocks by ergodic inclusions in the dynamics of the disordered spin chains. There are, however, two main caveats. The tendency towards ergodicity with increased time and length scales is similar to that of the uniformly disordered spin chains. This prevents distinguishing the dynamics of the MBL phase from a very slow thermalization. Moreover, the presence of two lengthscales $L_{A}$ and $L_{B}$ makes the extrapolation of the results to the thermodynamic limit even more complicated than in the case of the uniformly disordered XXZ spin chains. Nonetheless, the two-point correlation function (80) can be probed experimentally in many-body systems with planted thermal inclusions, as demonstrated in [393].
@@
Another approach to the investigation of the effects of planted regions of weak disorder was taken in [516], which considers the dynamics of the autocorrelation function $C(t)$ [^45] Strictly speaking, [516] considered the dynamics of the imbalance $I(t)$ from the initial Néel state. in the disordered XXZ spin chain with on-site potential drawn according to Eq. (78). The decay of $C(t)$ is strongly enhanced in the presence of thermal inclusion compared to the situation when $W_{A}=W_{B}$. Moreover, the results for $C(t)$ in the presence of the thermal inclusion collapse on a single curve when the time is rescaled according to $t\to te^{-aW}$, where $a=2.1$ is a constant. The rescaling shows a significant degree of uniformity of the dynamics across a broad regime of disorder strengths, suggesting thermalization of the system in the presence of thermal inclusion. However, the system considered in [516] is periodically driven by turning on and off the coupling $\hat{H}_{AB}$ between the ergodic inclusion and MBL region. The same system, but with coupling $H_{AB}$ present at all times, shows a much weaker tendency towards thermalization [685]. While the coupling to thermal inclusion indeed increases the tendency towards thermalization, the rescaling $t\to te^{-aW}$ does not provide a satisfactory collapse of the results for $C(t)$ in a broad regime of disorder strengths in the system without the periodic driving. This suggests disparate dynamics in the autonomous system at small and large $W_{A}$, leaving the question about its thermalization in the $L,t\to\infty$ limit open.
@@
The results of the studies of disordered many-body systems with planted thermal inclusions are consistent with the predictions of the avalanche scenario. However, the approach of [393, 685], based on investigation of the thermalization length $\xi_{d}(t)$, asserts the mere consistence of the thermalization in the subsystem $A$ with the propagation of avalanches. Therefore, it is important to note that the results of [685] do not exclude the possibility that the thermalization of the subsystem $A$ occurs due to microscopic mechanism different from the avalanche scenario, but with similar predictions for the behavior of $\xi_{d}(t)$. Whether such a mechanism could be formulated remains an open question. Nevertheless, for that reason, the problem of pin-pointing the microscopic mechanism responsible for the emergence of ergodicity in disordered spin chains remains a challenge for further investigations of the dynamics of disordered many-body systems.